OpenStudy (anonymous):

Write the following expression as a function of a positive acute angle. cos 125

terenzreignz (terenzreignz):

@aussy123 It might help you to know that with your angle in degrees, \[\large \cos \ \theta = \sin (90^o - \theta) \]

terenzreignz (terenzreignz):

So \[\large \cos(125^o)=\sin(90^o - 125^o)=\sin(-35^o)\]

OpenStudy (anonymous):

why did you use sin

terenzreignz (terenzreignz):

It's an identity. It didn't say to use a specific function, it just says use a function.

OpenStudy (anonymous):

oh sorry

terenzreignz (terenzreignz):

Get it now?

OpenStudy (anonymous):

yes a little

OpenStudy (anonymous):

the answer is -cos55

terenzreignz (terenzreignz):

I suppose so. You're right :) It stems from the fact that it's equal to sin(-35) = -sin(35) = -cos(90 - 35)

OpenStudy (anonymous):

I tried that way too, but I also did cos 125 = cos (180 - 125) = - cos 55

terenzreignz (terenzreignz):

Well, as long as your methods are valid, it should be no problem.

OpenStudy (anonymous):

Thank you!

terenzreignz (terenzreignz):

No problem :)